OpenStudy (sriram):

whats bigger e^pi or pi^e? do not use calculator and give an explanation for your ans.

OpenStudy (sriram):

hint: it can be solved using calculus

OpenStudy (anonymous):

\[(2.7)^{3.14}\] Or, \[(3.14)^{2.7}\]

OpenStudy (unklerhaukus):

the results are too close for a numerical approach

OpenStudy (sriram):

both e^pi and pi^e are positive now raise both to the power -e*pi which makes them e^1/e and pi^1/pi

OpenStudy (anonymous):

I think pi^e is greater...

OpenStudy (sriram):

now they are of the form x^1/x hence if u diff we get \[[x ^{1/x}/x ^{2}] * [1-\ln x]\]

OpenStudy (sriram):

which means f'x is negative for x >e hence as x increases after x=e f(x) keeps decreasing meaning pi^e < e^pi

OpenStudy (anonymous):

Here's one way: Consider the function f(x) = x / ln(x). Its derivative is f ' (x) = 1*(1/ln(x)) + x[ -(ln(x))^-2 ]*(1/x) f ' (x) = (1/ln(x)) - 1/(ln(x))^2 f ' (x) = (ln(x) - 1) / [ln(x)]^2 Setting this equal to 0 shows that it reaches a local minimum when x=e. So since e < pi, that must mean f(e) < f(pi). This means: e / ln(e) < pi / ln(pi) e < pi/ln(pi) e ln(pi) < pi ln(pi^e) < pi ln(pi^e) < pi * 1 ln(pi^e) < pi * ln(e) ln(pi^e) < ln(e^pi) pi^e < e^pi So e^pi is greater.

OpenStudy (sriram):

lovely man

OpenStudy (sriram):

hey nitz how did u think of it?